# Subsystem Thermalization Hypothesis in Quantum Spin Chains with Conserved Charges

### v1

@@ Abstract
We consider the thermalization hypothesis of pure states in quantum Ising chain with $Z_{2}$ symmetry, XXZ chain with $U(1)$ symmetry, and XXX chain with $SU(2)$ symmetries. Two kinds of pure states are considered: the energy eigenstates and the typical states evolved unitarily from the random product states for a long enough period. We further group the typical states by their expectation values of the conserved charges and consider the fine-grained thermalization hypothesis. We compare the locally (subsystem) reduced states of typical states/eigenstates with the ones of the corresponding thermal ensemble states. Besides the usual thermal ensembles such as the (micro-)canonical ensemble without conserved charges and the generalized Gibbs ensemble (GGE) with all conserved charges included, we also consider the so-called partial-GGEs (p-GGEs), which include only part of the conserved charges in the thermal ensemble. Moreover, in the framework of p-GGE, the Hamiltonian and other conserved charges are on an equal footing. The introduction of p-GGEs extends quantum thermalization to a more general scope. The validity of the subsystem thermalization hypothesis can be quantified by the smallness of the relative entropy of the reduced states obtained from the GGE/p-GGE and the typical states/eigenstates. We examine the validity of the thermalization hypothesis by numerically studying the relative entropy demographics. We show that the thermalization hypothesis holds generically for the small enough subsystems for various p-GGEs. Thus, our framework extends the universality of quantum thermalization.
@@ I Introduction
The thermalization of a classical many-body system is dictated by the second law of thermodynamics and is also deeply related to chaos and ergodicity. For quantum systems, a long-evolved open subsystem interacting with the environment displays similar thermalization behaviors, which can also be understood as the spreading of entanglement across the boundary between the subsystem and the environment. On the other hand, the isolated quantum system evolves unitarily, and it seems unable to thermalize. Among them, the energy eigenstate is stationary and should not be able to thermalize. Inspired by von Neuman’s idea of trying to generalize the ergodicity theorem to quantum mechanics [1], the so-called eigenstate thermalization hypothesis (ETH) was proposed [2, 3, 4], which states that the expectation value of any local observable with respect to an energy eigenstate of medium energy is approximately by the microcanonical ensemble average. Since then, thermalization of isolated quantum systems has been intensively studied in the past decades; see [5, 6, 7, 8, 9] and the references therein. Surprisingly, ETH seems quite general. The ETH implies that the reduced state in a small region of an energy eigenstate can be approximated by the (reduced) state of some thermal ensemble state, such as the (micro-)canonical ensemble. This can be seen as the subsystem version of ETH [10, 11, 12], i.e., by treating a local region as a subsystem of the whole system. The validity of (subsystem) ETH implies that an energy eigenstate locally looks like thermal.
@@
Energy eigenstates comprise only a tiny portion of the total Hilbert space of an isolated many-body system. It is natural to ask if the (subsystem) thermalization also happens to the typical states, which can be practically obtained by the long-time evolution of some initial non-eigenstate. This was proposed in [13] as the canonical typicality to conjecture that a typical state is locally thermal. The subsystem thermalization of a typical state of an isolated system may be understood by treating the subsystem as an open system surrounded by the rest. The open system will be expected to equilibrate with the surrounding environment due to their mutual interaction to grow and spread the entanglement across the boundary. Despite that, the equilibrium state may not be some thermal ensemble state, but some pointer state preferred by the environment’s einselection [14]. This is expected because the environment here is just complementary to the subsystem of a typical state and may not behave as a thermal bath. However, inspired by the subsystem ETH, the reduced state to the subsystem of a typical state can look like the reduced state of some thermal ensemble state [15, 16].
@@
It has been known that the above (subsystem) ETH fails [^1] Or, it holds only weakly (weak ETH) [17, 18, 19, 12], i.e., the distance between two reduced states vanishes by the power law of the number of degrees of freedom, not by the exponential decay as for the strong ETH. for the integrable systems [17, 18, 19, 12, 20, 21, 22] due to the infinite number of conserved charges prohibiting quantum chaos for thermalization. One expects the ETH will hold when compared with the reduced state of the generalized Gibbs ensemble (GGE), including all the conserved charges [23, 24]. However, such GGE is hard to implement due to the infinite number of conserved charges. Similarly, we may expect that the integrable systems’ typical states may not be thermalized if not compared with GGE. This then raises the issue of the role of the conserved charges in the thermalization of energy eigenstates or typical states.
@@
With the additional conserved charges other than Hamiltonian, there are more options when considering the thermalization hypothesis in various aspects. Firstly, the energy eigenstates will be classified into the superselection sectors of the conserved charges. When considering ETH, we can choose to restrict to the superselection sectors or not. We can also generalize the concept of superselection sectors to the typical states. Although the typical states, in general, are not the eigenstates of energy and charges, we can still classify them into almost superselection by their expectation values of the charges.
@@
Secondly, we can choose different thermal ensemble states to examine the thermalization hypothesis for either eigenstates or typical states. A conserved quantity and its chemical potential come as a conjugate pair when constructing a thermal ensemble, i.e., we can fix either charge or chemical potential, but not both. Or, we can ignore both. For example, when there is no conserved quantity other than the Hamiltonian, we can either fix energy to yield a microcanonical ensemble state or fix (inverse) temperature to yield a canonical ensemble state. We can adopt either to consider the thermalization hypothesis. Thus, with more conserved quantities, we can fix either one for each conjugate pair or ignore both to yield different thermal ensembles. If we fix all the chemical potential, including the inverse temperature, we will obtain the GGE. Otherwise, we will obtain the so-called partial-GGE (p-GGE). Finally, as we have the option of choosing (almost) superselection sectors for the quantum states and the p-GGEs, we should not just examine the thermalization of a single state but shall evaluate the thermalization hypothesis by the Demographics of the relative entropy measures for a given set of states in some (almost) superselection sector and a chosen p-GGE.
@@
In this paper, we will study this issue by considering the nonintegrable spin chain systems with either discrete or continuous Abelian and non-Abelian symmetries. In particular, the thermalization hypothesis for the system with non-Abelian conserved charges has been studied recently [25, 26, 27, 28, 29, 30], among which a GGE-like state called non-Abelian thermal state (NATS) has been proposed for ETH. In our framework with p-GGE states, we will numerically study the subsystem thermalization hypothesis of the typical states, (almost-)superselection states, and energy eigenstates of the quantum Ising and XXZ chains by comparing their locally reduced states with the ones of various p-GGE thermal states. The demographics of such comparisons will tell us how general the thermalization of isolated quantum systems with conserved charges can be.
@@
The rest of the paper is organized as follows. In section II, we will sketch our framework for examining the subsystem thermalization hypothesis and its relation with the generalized second law of thermodynamics and our numerical implementation methodology. We then present our numerical results of the thermalization demographics for the Ising chain with $Z_{2}$ conserved charges, the XXZ chain with $U(1)$ charges, and the XXX chain with $SU(2)$ charges, in section III, IV and V, respectively. We will see that the subsystem thermalization hypothesis holds well for the local region of small sizes in most cases. Finally, we conclude our paper in section VI. To make the main text more concise, we put some typical-state thermalization demographics for the cases of XXZ chain in Appendix A.
@@ II Thermalization of Quantum Systems with Conserved Charges
There are many ways to consider the thermalization of a quantum system, and the key point is to pick up a thermal ensemble state to compare with the given typical state on the agreement of local observables or, more directly, the reduced density matrices of some local regions. Take the eigenstate thermalization hypothesis (ETH) [2, 4] as an example. One compares a given energy eigenstate $|E\rangle$ of energy E with some thermal ensemble state, which can either be a micro-canonical or canonical ensemble. It is natural to choose the energy of the micro-canonical ensemble to center around $E$ with a small window $\Delta E\leq E$. The ETH works if the local observables’ expectation values agree well with the micro-canonical ensemble’s thermal averages. That is
$\langle E|{\cal O}|E\rangle\simeq\langle{\cal O}\rangle_{\rm mce}$ (1)
for a complete set of local observables $\{{\cal O}\}$, where ${\rm mce}$ denotes the abbreviation of micro-canonical ensemble. On the other hand, the canonical ensemble
@@
is specified by the system’s Hamiltonian $H$ and the inverse temperature $\beta$. The latter should be fixed by requiring the average energy to be $E$, i.e.,
@@
The ETH holds if the reduced density matrices for a small local region $A$ (with its complement denoted by $\overline{A}$) agree, i.e.,
@@
where $S\big{[}\rho_{1}\|\rho_{2}\big{]}$ is the relative entropy between $\rho_{1}$ and $\rho_{2}$. This is the so-called subsystem ETH [10, 11, 12]. A tighter measure than the relative entropy for the difference between $\rho_{1}$ and $\rho_{2}$ is the trace distance $t_{12}:={\rm tr}|\rho_{1}-\rho_{2}|$. However, it is more difficult to evaluate. Despite that, by Pinsker’s inequality
@@
Besides ETH, one can consider the thermalization for a typical pure state $|\Psi\rangle$, which is not an eigenstate or a product state. Then, the canonical ensemble or its $\beta$ can be fixed by
@@
In [31], the above thermalization hypothesis is proposed and termed generic ETH, which has been checked for the long-time evolved product states. In this work, we call it the subsystem thermalization hypothesis, a generalization of subsystem ETH to include both eigenstates and typical states. For the subsystem thermalization hypothesis to hold, the size of $A$ should be far smaller than the system’s. As $A$ increases its size to some critical one denoted by $A_{c}$, one should expect the failure of the subsystem thermalization hypothesis. Thus, $A_{c}$ can be used to characterize the robustness of the thermalization hypothesis and should be system-dependent. For example, one will wonder how the symmetries of the system affect $A_{c}$.
@@
We can generalize the above subsystem thermalization hypothesis to systems with conserved charges $\{Q_{I}\}$ where $I$ labels the charges. For simplicity, we set $Q_{0}=H$ so that $[Q_{0},Q_{I\neq 0}]=0$. In statistical mechanics, the conserved charge $Q_{I}$ and its corresponding chemical potential $\beta_{I}$ form the conjugate pair. When fixing a thermal ensemble, we can fix either $Q_{I}$ or $\beta_{I}$. Denote the set of fixed charges as $\{Q_{a}\}$ and the set of fixed chemical potentials as $\{\beta_{j}\}$. These two sets are mutually exclusive, i.e., $\{a\}\cap\{j\}=\emptyset$. However, there are remaining charges labeled by index $\gamma$, for which we fix neither $Q_{\gamma}$ nor $\beta_{\gamma}$. Note that $\{a\}\cup\{j\}\cup\{\gamma\}=\{I\}$. With the above classification of the conserved charges, we define the so-called partial generalized Gibbs ensemble (p-GGE) thermal state as
@@
After introducing the p-GGE, the subsystem thermalization hypothesis can be extended to the systems with conserved charges. For a given state $|\Psi\rangle$, a chosen p-GGE ensemble state can be fixed by requiring
@@
    Extended ETH. In this case, we choose $\{Q_{a}\}$ to be mutually commuting and examine the thermalization hypothesis of their eigenstates |$|\{q_{a}\}\rangle$, defined by
@@
    for a given set $\{\langle Q_{a}\rangle\}$. This case is the fine-grained version of the typical-state thermalization hypothesis by subdividing the typical states into almost superselection sectors specified by $\langle Q_{a}\rangle$.
@@
    When considering the subsystem thermalization hypothesis for the non-Abelian symmetry systems, we can also choose the almost superselection states $\{|\Psi^{\{a\}}\rangle\}$ for the non-commuting $\{Q_{a}\}$, or the eigenstates {|$\{|\{q_{a}\}\rangle\}$ for commuting $\{Q_{a}\}$. These provide new ways to examine the thermalization hypothesis of typical states and ETH with non-Abelian symmetries.
@@
The main goal of this paper is to study the thermalization demographic for each possible scenario discussed above and see if the thermalization is generic enough. The relative entropy is adopted to characterize the thermalization demographics, and its notations used later for various scenarios are summarized in Table 1.
In summary, we will consider the subsystem thermalization hypothesis for two types of spin-$1/2$ chains with conserved charges: the quantum Ising chains and the quantum XXZ chains. The nonintegrable Ising chains have a conserved $Z_{2}$ charge. Interestingly, as we will see, the subsystem thermalization hypothesis also works for the p-GGE of discrete charge. The symmetries of XXZ chains are richer than the Ising’s, but we will mainly focus on $U(1)$ and $SU(2)$. The latter is non-Abelian, and it is interesting to examine the subsystem thermalization hypothesis for typical states or eigenstates of systems with such non-Abelian symmetries by comparing them with p-GGE.
@@
Table 1: Notations of realtive entropy $S_{A}$ defined in (10) and used in different scenarios of the thermalization hypothesis. The superscript $t$ denotes the typical state, $e$ is the eigenstate, and $a$ is the almost superselection or just superselection state. The size of $A$ ranges from one site to half of the spin chain.
Before presenting the thermalization demographics, we have two remarks about the generality of quantum thermalization and how to prepare the typical states numerically. Firstly, we consider the generality of quantum thermalization by extending an argument for GGE in [32] to the cases for p-GGE. This can provide some understanding of quantum thermalization by the second-law-like argument. The key idea of [32] is to define the so-called free entropy for a generic state $\rho$ for a given set of chemical potentials $\{\beta_{j}\}$ as follows [32],
@@
This is the statement of the second law of thermodynamics. Given free entropy as the monotone for state transitions, the free operations are just the evolution conserving $Q_{j}$’s, e.g., $U(t)=e^{-it\sum_{j}\alpha_{j}Q_{j}}$, which will not change the free entropy. Otherwise, the operations are the so-called thermal operations (TOs) [33].
@@
Secondly, we will mention how we numerically prepare and define the typical states. The subsystem thermalization hypothesis can work for the systems even without taking the thermodynamic limit as long as the subsystem size is far smaller than the system’s. We will mainly adopt the exact diagonalization method (ED) to find the eigenstates of the spin chain models with the size of about ten sites. We will also use it to evolve the initial product states to prepare the typical states for checking the subsystem thermalization hypothesis. Explicitly, we first prepare the initial entangled state by $|\psi\rangle=\otimes_{i=1}^{L}\Big{(}\cos(\theta_{i}/2)|0\rangle+e^{i\phi_{i}}\sin(\theta_{i}/2)|1\rangle\Big{)}$ with $\theta_{i}\in[0,\pi]$, $\phi_{i}\in[0,2\pi]$ generated uniformly and randomly. Then, we evolve these states with the Hamiltonian of the spin chain model by ED for chains of 10 sites or so. Although 10-site chains seem small, as we will see, it is enough to achieve the subsystem thermalization at least for $A=1$. One will expect $A_{c}$ to increase as the system size increases. This can be a future study by using matrix produce states (MPS) to obtain the typical states for the longer chains. In our previous work [34], we have shown it is possible to accurately evolve product states or prepare thermal states using matrix product states (MPS) for spin chains of a few hundred sites.

### v2

@@ Abstract
We consider the thermalization hypothesis of pure states in non-integrable quantum spin chains with conserved charges, including the non-intergable variants of the Ising chain with $Z_{2}$ symmetry, XXZ chain with $U(1)$ symmetry, and XXX chain with $SU(2)$ symmetries. Two kinds of pure states are considered: the energy eigenstates and the typical states evolved unitarily from the random product states for a long enough period. We further group the typical states by their expectation values of the conserved charges and consider the fine-grained thermalization hypothesis. We compare the locally (subsystem) reduced states of typical states/eigenstates with the ones of the corresponding thermal ensemble states. Besides the usual thermal ensembles such as the (micro-)canonical ensemble without conserved charges and the generalized Gibbs ensemble (GGE) with all conserved charges included, we also consider the so-called partial-GGEs (p-GGEs), which include only part of the conserved charges in the thermal ensemble. Moreover, in the framework of p-GGE, the Hamiltonian and other conserved charges are on an equal footing. The introduction of p-GGEs extends quantum thermalization to a more general scope. The validity of the subsystem thermalization hypothesis can be quantified by the smallness of the relative entropy of the reduced states obtained from the GGE/p-GGE and the typical states/eigenstates. We examine the validity of the thermalization hypothesis by numerically studying the relative entropy demographics. We show that the thermalization hypothesis holds generically for small enough subsystems for various p-GGEs. Thus, our framework extends the universality of quantum thermalization.
@@ I Introduction
The thermalization of a classical many-body system is governed by the second law of thermodynamics and is also closely linked to chaos and ergodicity. For quantum systems, a long-evolved open subsystem interacting with the environment displays similar thermalization behaviors, which can also be understood as the spreading of entanglement across the boundary between the subsystem and the environment. On the other hand, the isolated quantum system evolves unitarily and appears unable to thermalize. Among them, the energy eigenstate is stationary and should not be able to thermalize. Inspired by von Neuman’s idea of trying to generalize the ergodicity theorem to quantum mechanics [1], the so-called eigenstate thermalization hypothesis (ETH) was proposed [2, 3, 4, 5], which states that, even in isolated quantum system, the expectation value of a local observable in a single energy eigenstate is well approxemated by the microcanonical ensemble average with finite energy density, up to corrections that vanish in the thermodynamic limit. Since then, the thermalization of isolated quantum systems has been intensively studied over the past few decades; see [6, 7, 8, 9, 10] and the references therein. Surprisingly, ETH seems quite general. The ETH implies that the reduced state in a small region of an energy eigenstate can be approximated by the (reduced) state of some thermal ensemble state, such as the (micro-)canonical ensemble. This can be viewed as the subsystem version of ETH [11, 12, 13], i.e., by treating a local region as a subsystem of the entire system. The validity of (subsystem) ETH implies that an energy eigenstate locally looks like thermal.
@@
Energy eigenstates comprise only a tiny portion of the total Hilbert space of an isolated many-body system. It is natural to ask if the (subsystem) thermalization also happens to the typical states, which can be practically obtained by the long-time evolution of some initial non-eigenstate. This was proposed in [14] as the canonical typicality to conjecture that a typical state is locally thermal. [^1] The typical states adopted in [14, 15, 16] for non-integrable systems are the Haar-random states from the system’s Hilbert (sub-)space. However, due to the limited computing power and the purpose of this paper, we will prepare the typical states by evolving the product of on-site random states, which sample the system’s Hilbert space well, for a sufficiently long time to saturate the local entanglement. This should be sufficient for considering the thermalization of local observables due to the similar mechanism of scrambling by unitary evolution proposed in [17, 18], with supporting examples as demonstrated in [6, 19, 20]. The subsystem thermalization of a typical state of an isolated system may be understood by treating the subsystem as an open system surrounded by the rest. The open system is expected to equilibrate with the surrounding environment due to their mutual interaction, allowing the entanglement to grow and spread across the boundary. Despite that, the equilibrium state may not be some thermal ensemble state, but some pointer state preferred by the environment’s einselection [21]. This is expected because the environment here is just complementary to the subsystem of a typical state and may not behave as a thermal bath. However, inspired by the subsystem ETH, the reduced state of a subsystem in a typical state can resemble the reduced state of some thermal ensemble state [15, 16].
@@
It has been known that the above (subsystem) ETH fails [^2] Or, it holds only weakly (weak ETH) [22, 23, 24, 13], i.e., the distance between two reduced states vanishes by the power law of the number of degrees of freedom, not by the exponential decay as for the strong ETH. Besides, there exist atypical states, which comprise an exponentially small fraction in system size and violate weak ETH [23]. for the integrable systems [22, 23, 24, 13, 25, 26, 27, 28, 29] due to the infinite number of conserved charges prohibiting quantum chaos for thermalization. It also fails for many-body localized systems [30, 31, 32, 33] and chaotic systems in the presence of scars [34, 35, 36, 37, 38, 39], with different underlying mechanisms, such as strong disorders for the former, or Hilbert space fragmentation due to emerging spectrum generating algebra for the latter. One expects the ETH to hold for the generalized Gibbs ensemble (GGE), which includes all conserved charges [40, 41]. However, such GGE states are hard to implement due to the infinite number of conserved charges, and one may want to adopt some truncated GGE states. Indeed, some studies on the approximation of GGE states by the truncated GGEs have been done in [42, 43, 44] by including some semi-local charges or imposing integrable initial conditions. Similarly, we may expect the integrable systems’ typical states also satisfy the subsystem thermalization hypothesis with GGE or some truncated GGE. This then raises the issue of the role of the conserved charges in the thermalization of energy eigenstates or typical states.
@@
With the additional conserved charges other than Hamiltonian, there are more options when considering the thermalization hypothesis in various aspects. Firstly, the energy eigenstates will be classified into the superselection sectors of the conserved charges. When considering ETH, we can choose to restrict to the superselection sectors or not. We can also generalize the concept of superselection sectors to the typical states. Although typical states, in general, are not eigenstates of energy and charges, we can still classify them into almost superselection sectors by their expectation values of the charges.
@@
Secondly, we can choose different thermal ensemble states to examine the thermalization hypothesis for either eigenstates or typical states. A conserved quantity and its chemical potential form a conjugate pair when constructing a thermal ensemble; i.e., we can fix either charge or chemical potential, but not both. Or, we can ignore both. For example, when there is no conserved quantity other than the Hamiltonian, we can either fix energy to yield a microcanonical ensemble state or fix (inverse) temperature to yield a canonical ensemble state. We can adopt either to consider the thermalization hypothesis. Thus, with more conserved quantities, we can fix either one for each conjugate pair or ignore both to yield different thermal ensembles. If we fix all the chemical potentials, including the inverse temperature, we will obtain the GGE. Otherwise, we will obtain the so-called partial-GGE (p-GGE). The freedom to choose GGE or various p-GGE implies that one can choose to ignore specific information when considering the thermalization hypothesis. As most of the studies on the thermalization hypothesis are based on GGE [40, 41, 45, 46, 47, 48, 49, 50, 51], we propose in this paper to study the thermalization hypothesis of typical states and energy eigenstates by comparing them with various p-GGE ensemble states. Our results indicate that the thermalization hypothesis is effective in most cases. This implies that the subsystem thermalization hypothesis can be extended from GGE to p-GGEs. This then enlarges the scope of quantum thermalizations for the systems with conserved charges.
Finally, as we have the option of choosing (almost) superselection sectors for the quantum states and the p-GGEs, we should not just examine the thermalization of a single state but shall evaluate the thermalization hypothesis by the demographics of the relative entropy measures for a given set of states in some (almost) superselection sector and a chosen p-GGE. Here, we adopt the term ”demographic” to mean the population distribution of the chosen ensemble typical states/eigenstates over the values of the relative entropy. If the demographic mostly peaks around the regime of vanishing relative entropy, the thermalization hypothesis works for the chosen ensemble states and p-GGE.
In particular, experiments such as those by Kaufman et al.[52] have demonstrated the applicability of the quantum thermalization framework in cold atom and Rydberg atom platforms, where coherent many-body dynamics and local observables can be directly probed. Other platforms, including trapped ion systems[53] and superconducting qubit arrays [54], also provide a high degree of control over isolated quantum many-body systems. These setups enable the preparation of specific initial states, coherent unitary evolution, and local measurements, providing an environment for investigating thermalization phenomena.
In parallel, recent experimental advances have made it possible to explore quantum many-body systems with multiple conserved quantities or near-integrable dynamics. For example, Kinoshita, Wenger, and Weiss [55] observed the absence of thermalization in a 1D Bose gas, a hallmark of integrability often referred to as a ”quantum Newton’s cradle.” Gring et al.[56] studied prethermalization in isolated quantum systems, while Langen et al.[57] provided experimental evidence for the emergence of a generalized Gibbs ensemble (GGE), which incorporates multiple conserved quantities beyond energy. More recently, Scholl et al. [58] used programmable Rydberg atom arrays to simulate two-dimensional antiferromagnetic spin models, enabling the study of constrained dynamics and engineered local symmetries. These results highlight that nontrivial conserved structures are not only theoretically relevant but also experimentally realizable, thereby motivating further investigations into their role in quantum thermalization.
@@
In this paper, we will study quantum thermalization in the above setup by considering the non-integrable spin chain systems with either discrete or continuous Abelian and non-Abelian symmetries. In particular, the thermalization hypothesis for systems with non-Abelian conserved charges has been studied recently [45, 46, 47, 48, 49, 50, 51], among which a GGE-like state called non-Abelian thermal state (NATS) has been proposed for ETH. In our framework with p-GGE states, we will numerically study the subsystem thermalization hypothesis for the typical states, (almost-)superselection states, and energy eigenstates of non-integrable variants of the quantum Ising and XXZ chains by comparing their locally reduced states with those of various p-GGE thermal states. The demographics of such comparisons will reveal how general the thermalization of isolated quantum systems with conserved charges is.
@@
The rest of the paper is organized as follows. In Section II, we will outline our framework for examining the subsystem thermalization hypothesis and its relationship with the generalized second law of thermodynamics, as well as our numerical implementation methodology. We then present our numerical results of the thermalization demographics for the (non-integrable variants of) Ising chain with $Z_{2}$ conserved charges, the XXZ chain with $U(1)$ charges, and the XXX chain with $SU(2)$ charges, in Section III, IV and V, respectively. We will see that the subsystem thermalization hypothesis holds well for the local region of small sizes in most cases. Finally, we conclude our paper in Section VI. In Appendix A, we conduct a size-scaling study of the thermalization demographics of Ising chains by varying the system size. To make the main text more concise, we put some typical-state thermalization demographics for the cases of XXZ chains in Appendix B with $U(1)$ charge and in Appendix C with parity symmetry.
@@ II Thermalization of Quantum Systems with Conserved Charges
There are many ways to consider the thermalization of a quantum system, and the key point is to pick up a thermal ensemble state to compare with the given typical state on the agreement of local observables or, more directly, the reduced density matrices of some local regions. Take the eigenstate thermalization hypothesis (ETH) [2, 4, 5] as an example. One compares a given energy eigenstate $|E\rangle$ of energy $E$ with a thermal ensemble state, which can be either a microcanonical or canonical ensemble. It is natural to choose the energy of the micro-canonical ensemble to center around $E$ with a small window $\Delta E\leq E$. The ETH works if both diagonal and off-diagonal matrix elements of a set of local operators $\mathcal{O}$ in the basis of eigenstates satisfy the following Srednicki’s ansatz [5],
$\langle E_{\alpha}|\mathcal{O}|E_{\beta}\rangle\simeq\langle\mathcal{O\rangle_{\rm mce}}(\overline{E})\delta_{\alpha\beta}+e^{-S_{\rm th}(\overline{E})/2}R_{\alpha\beta}f(\omega,\overline{E}),$ (1)
where $\rm mce$} denotes the abbreviation of micro-canonical ensemble, $\overline{E}=(E_{\alpha}+E_{\beta})/2$ denotes the average eigenenergy, $\omega=E_{\alpha}-E_{\beta}$ is the energy difference, $S_{\rm th}(\overline{E})$ is the thermodynamics entropy, $R_{\alpha\beta}$ is a normal distributed random number, and $f(\omega,\overline{E})$ is a smooth spectral function.
On the other hand, we can consider the so-called subsystem ETH [11, 12, 13] based on the canonical ensemble
@@
which is specified by the system’s Hamiltonian $H$ and the inverse temperature $\beta$. The latter should be fixed by requiring the average energy to be the energy $E$ of the energy eigenstate $|E\rangle$, i.e.,
@@
Unlike the microcanonical ensemble, the canonical ensemble allows energy fluctuations but fixes the inverse temperature. The subsystem ETH holds if the reduced density matrices for a small local region $A$ (with its complement denoted by $\overline{A}$) agree, i.e.,
@@
where $S\big{[}\rho_{1}\|\rho_{2}\big{]}$ is the relative entropy between $\rho_{1}$ and $\rho_{2}$. A tighter measure than the relative entropy for the difference between $\rho_{1}$ and $\rho_{2}$ is the trace distance $t_{12}:={\rm tr}|\rho_{1}-\rho_{2}|$. However, it is more difficult to evaluate. Despite that, by Pinsker’s inequality [59]
@@
Besides (subsystem) ETH, one can consider the thermalization for a typical pure state $|\Psi\rangle$, which is not an eigenstate or a product state. Then, the canonical ensemble or its $\beta$ can be fixed by
@@
In [60], the above thermalization hypothesis is proposed and termed the generic ETH, which has been verified for long-time evolved product states. In this work, we refer to it as the subsystem thermalization hypothesis, a generalization of the subsystem ETH to include both eigenstates and typical states. For the subsystem thermalization hypothesis to hold, the size of $A$ should be significantly smaller than that of the system. In Appendix A, we conduct a limited size-scaling study to provide evidence that the thermalization hypothesis is more effective for larger systems.
@@
We can generalize the above subsystem thermalization hypothesis to systems with conserved charges $\{Q_{I}\}$ where $I$ labels the charges. For simplicity, we set $Q_{0}=H$ so that $Q_{0},Q_{I\neq 0}]=0$. In statistical mechanics, the conserved charge $Q_{I}$ and its corresponding chemical potential $\beta_{I}$ form the conjugate pair. When fixing a thermal ensemble, we can fix either $Q_{I}$ or $\beta_{I}$. Denote the set of fixed charges as $\{Q_{a}\}$ and the set of fixed chemical potentials as $\{\beta_{j}\}$. These two sets are mutually exclusive, i.e., $\{a\}\cap\{j\}=\emptyset$. However, there are remaining charges labeled by index $\gamma$, for which we fix neither $Q_{\gamma}$ nor $\beta_{\gamma}$. Note that $\{a\}\cup\{j\}\cup\{\gamma\}=\{I\}$. With the above classification of the conserved charges, we define the so-called partial generalized Gibbs ensemble (p-GGE) thermal state as
@@
After introducing the p-GGE, the subsystem thermalization hypothesis can be extended to systems with conserved charges. For a given state $|\Psi\rangle$, a chosen p-GGE ensemble state can be fixed by requiring
@@
    Extended ETH. In this case, we choose $\{Q_{a}\}$ to be mutually commuting and examine the thermalization hypothesis of their eigenstates $|\{q_{a}\}\rangle$, defined by
@@
    for a given set $\{\langle Q_{a}\rangle\}$. In our work, we define the ”almost superselection states” by the condition that the difference between the corresponding quantities is approximately $0.1\%$ of $\langle Q_{a}\rangle$, which we consider to be a reasonably small and acceptable uncertainty for our analysis. This case is the fine-grained version of the typical-state thermalization hypothesis, which subdivides the typical states into almost superselection sectors specified by $\langle Q_{a}\rangle$.
@@
    When considering the subsystem thermalization hypothesis for the non-Abelian symmetry systems, we can also choose the almost superselection states $\{|\Psi^{\{a\}}\rangle\}$ for the non-commuting $\{Q_{a}\}$, or the eigenstates $\{|\{q_{a}\}\rangle\}$ for commuting $\{Q_{a}\}$. These provide new ways to examine the thermalization hypothesis of typical states and ETH with non-Abelian symmetries.
@@
The primary objective of this paper is to investigate the thermalization demographics for each possible scenario discussed above and determine if the thermalization is sufficiently generic. As mentioned, we will adopt the ensemble states’ demographics over the relative entropy to characterize the validity of the thermalization hypothesis for the corresponding p-GGE. The associated notations used later for various scenarios are summarized in Table 1.
In summary, we will consider the subsystem thermalization hypothesis for two types of spin-$1/2$ chains with conserved charges: the quantum Ising chains and the quantum XXZ chains. The nonintegrable Ising chains have a conserved $Z_{2}$ charge. Interestingly, as we will see, the subsystem thermalization hypothesis also works for the p-GGE of discrete charge. The symmetries of XXZ chains are richer than Ising’s, but we will mainly focus on $U(1)$ and $SU(2)$. The latter is non-Abelian, and it is interesting to examine the subsystem thermalization hypothesis for typical states or eigenstates of systems with such non-Abelian symmetries by comparing them with p-GGE.
@@
Table 1: Notations of realtive entropy $S_{A}$ defined in (10) and used in different scenarios of the thermalization hypothesis. The superscript $t$ denotes the typical state, $e$ is the eigenstate, and $a$ is the almost superselection or just superselection state. The size of AA ranges from one site to half of the spin chain.
Before presenting the thermalization demographics, we have two remarks about the generality of quantum thermalization and how to prepare the typical states numerically. Firstly, we consider the generality of quantum thermalization by extending an argument for GGE in [61] to the cases for p-GGE. This can provide some insight into quantum thermalization through a second-law-like argument. The key idea of [61] is to define the so-called free entropy for a generic state $\rho$ for a given set of chemical potentials $\{\beta_{j}\}$ as follows [61],
@@
This is the statement of the second law of thermodynamics. Given free entropy as the monotone for state transitions, the free operations are just the evolution conserving $Q_{j}$’s, e.g., $U(t)=e^{-it\sum_{j}\alpha_{j}Q_{j}}$, which will not change the free entropy. Otherwise, the operations are referred to as thermal operations (TOs) [62].
@@
Secondly, we will mention how we numerically prepare and define the typical states. The subsystem thermalization hypothesis can be applied to systems even without taking the thermodynamic limit, as long as the subsystem size is significantly smaller than that of the system. We will primarily employ the exact diagonalization method (ED) to determine the eigenstates of spin chain models with approximately ten sites. We will also use it to evolve the initial product states, preparing the typical states for checking the subsystem thermalization hypothesis. Explicitly, we first prepare the initial entangled state by $|\psi\rangle=\otimes_{i=1}^{L}\Big{(}\cos(\theta_{i}/2)|0\rangle+e^{i\phi_{i}}\sin(\theta_{i}/2)|1\rangle\Big{)}$ with $\theta_{i}\in[0,\pi]$, $\phi_{i}\in[0,2\pi]$ generated uniformly and randomly. Then, we evolve these states with the Hamiltonian of the spin chain model by ED for chains of 10 sites or so. If the evolution is long enough, we will expect the final states to be the typical states, leading to the thermalization of local observables due to the scrambling by the long-time unitary evolution [17, 18]. This can be checked by examining the saturation of the entanglement entropy. Although 10-site chains seem small, as we will see, they are sufficient to achieve subsystem thermalization, at least for $A=1$. One would expect $A_{c}$ to increase as the system size increases. This could be a future study using matrix product states (MPS) to obtain the typical states for longer chains. In our previous work [63], we demonstrated that it is possible to accurately evolve product states or prepare thermal states using matrix product states (MPS) for spin chains with a few hundred sites.